OpenStudy (anonymous):

Using log8 10=P , log8 12=Q, log8 11=R Find log8 5/256

OpenStudy (lgbasallote):

do you mean express it in letters?

OpenStudy (anonymous):

start with \[\log_8(5)-\log_8(256)\]

OpenStudy (anonymous):

Yes >.<

OpenStudy (anonymous):

then we see if we can write 5 and 256 with the numbers given.

OpenStudy (lgbasallote):

i think we're gonna need to add a few lol

OpenStudy (anonymous):

damn i got stuck a second hold on we know \[256=2^8\] so using base 8 we can write it as \[8^{\frac{8}{3}}\] so we know \[\log_8(256)=\frac{8}{3}\]

OpenStudy (anonymous):

Its hard huh ? I know the answer but i want to know how to get there.

OpenStudy (anonymous):

and since \(5=\frac{10}{2}\) we have \[\log_8(10)=\log_8(5)-\log_8(2)\]

OpenStudy (lgbasallote):

\[\log_8 5 = \log_8 10 - \log_8 2\]

OpenStudy (anonymous):

\[\log_8(2)=\frac{1}{3}\] because \[8^{\frac{1}{3}}=2\]

OpenStudy (anonymous):

It's P - 3 , some how.

OpenStudy (anonymous):

and you are given that \[\log_8(5)=P\]so your final answer is \[P-\frac{1}{3}-\frac{8}{3}=P-3\]

OpenStudy (anonymous):

i sense that this was not clear, so should we run through the steps again?

OpenStudy (anonymous):

Answer is P-3 . but P = log8 10

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

right so lets make sure it is clear what happened ok?

OpenStudy (anonymous):

okay (:

OpenStudy (anonymous):

\[\log_8(5)=\log_8(\frac{10}{2})=\log_8(10)-\log_8(2)\]

OpenStudy (anonymous):

why did we write it this way? because you were given that \[\log_8(10)=P\] so we could use that

OpenStudy (anonymous):

then since we can write \[8^{\frac{1}{3}}=2\] we know \[\log_8(2)=\frac{1}{3}\]

OpenStudy (anonymous):

And your using one of the properties of logarithms in which log x/y=log x-log y

OpenStudy (anonymous):

therefore we know \[\log_8(5)=P-\frac{1}{3}\] yes that is what i am using

OpenStudy (anonymous):

i also used it at the bigginning when i wrote \[\log_8(\frac{5}{256})=\log_8(5)-\log_8(256)\]

OpenStudy (lgbasallote):

@satellite73 got drunk a while ago and wrote \[\large \log_8 10 = \log_8 5 - \log_8 2\] lol

OpenStudy (anonymous):

yes you are right (on both counts) where did you get the bike??

OpenStudy (lgbasallote):

idk...i just googled red bike haha =))

OpenStudy (anonymous):

Try this one :) \[\log _{7}3 = X , \log _{7}8 = Y , \log _{7}10=Z Find \log _{7} 15/32\]